The table below lists the price and consumption levels of food and clothing during 2010 and 2020. Calculate a Laspeyres and Paasche index using 2010 as the base year (round your answer to three decimal points). [4] Year Prices Consumption Food Clothing Food Clothing 2010 500 300 200 85 2020 625 335 210 97


1
Expert's answer
2022-03-01T10:53:27-0500

Laspeyres="\frac{\Sigma P_n Q_0}{\Sigma P_0Q_0}\times 100"

= "\frac{(625\times 200)+(300\times85)}{(500\times 200)+ 335\times 85)}= \frac{78625}{128475} \times 100" = 61.20%


Paasche"= \frac{\Sigma P_n Q_n}{\Sigma P_0Q_n}\times 100"

"\frac{(625\times 210)+(300\times97)}{(500\times 210)+ 335\times 97)}= \frac{160350}{137495} \times 100" = 116.59%